OpenStudy (onew12):

Solve the following system of equations. 2x + y = 3 x = 2y - 1 I think {(1,1)} <-- this is the answer. Is that correct??

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 2x + y = 3 \quad \Longrightarrow\quad2\color{#ff0000}{\left(2y-1\right)}+y=3 }\) \(\color{#000000 }{ \displaystyle x = 2y - 1}\)

OpenStudy (solomonzelman):

I am just showing you that given the second equation, you can substitute 2y-1 instead of x, into the first equation.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 2\color{#ff0000}{\left(2y-1\right)}+y=3 }\) Can you solve for y?

OpenStudy (dayakar):

your answer is correct . show your work

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is an easy check, put x = 1, y = 1 and see if they both work (they do)

OpenStudy (solomonzelman):

Oh, lol, I didn't see the answer (just woke up).... I apologize:)

OpenStudy (onew12):

It's Ok :)